OpenStudy (anonymous):

How many we can arrange the word SESI? How many we can arrange the word SESE? How many we can arrange the word SERI?

OpenStudy (anonymous):

24 ways

OpenStudy (anonymous):

There are three different questions so there should be three different answers.

OpenStudy (anonymous):

in the first case, you have 4 letters you are rearranging, so thats 4! = 24, but two of the letters are the same. So you divide by 2! = 2 So you final answer is 24/2 = 12.

OpenStudy (anonymous):

That should tell you how to do the rest. Its always going to be 4!, but you divide if any of the letters are repeated.

OpenStudy (amistre64):

3 different questions which all rely on the same logic to solve ...

OpenStudy (amistre64):

the process is more important than the answers :)

OpenStudy (amistre64):

when we have multiple terms that are identical; the formula is: \[\frac{n!}{n_1!n_2!n_3!...n_k!}\] if i recall correctly

OpenStudy (anonymous):

right right :)

OpenStudy (anonymous):

How many we can arrange the word SESI? For this question 72 is the answer..

OpenStudy (anonymous):

(4!/2!) * 6

OpenStudy (anonymous):

Can anyone explain this answer?

OpenStudy (anonymous):

There arent even 72 ways to arrange 4 things, there are 4! = 24. No way the answer can be bigger than 24.

OpenStudy (anonymous):

where did that times 6 come from?

OpenStudy (anonymous):

4!/2! is the ways of arranging SESI .but among them we also need to arrange. Taking SS as one and E and I as other two.Among themselves we can arrange in 3! ways. Thus total no of ways (4!/2!) * 3!

OpenStudy (amistre64):

\begin{array}c &&s&i&-sesi\\ &e\\ &&i&s&-seis\\ &&e&i&-ssei\\ s&s\\ &&i&e&-ssie\\ &&e&s&-sies\\ &i\\ &&s&e&-sise\\ &&s&i\\ &s\\ &&i&s\\ &&s&i\\ e&s\\ &&i&s\\ &&s&s\\ &i\\ &&s&s\\ &&e&i\\ &s\\ &&i&e\\ &&s&i\\ s&e\\ &&i&s\\ &&s&e\\ &i\\ &&e&s\\ &&s&e\\ &s\\ &&e&s\\ &&s&s\\ i&e\\ &&s&s\\ &&e&s\\ &s\\ &&s&e\\ \end{array} etc...